You volunteered to be part of the medical mission in your barangay and in-charge to get weight of the children. The weights of 1,000 children have an average of 45 kilograms and a standard deviation of 5 kilograms.





a. How many children weigh 35 kg to 55 kg?

1
Expert's answer
2022-03-15T16:29:09-0400

Let X be a random variable representing the weight of the children, then X ~ "N(45,5^2)"

"P(35<X<55)=P(35<N(45,5^2)<55)=P(35<45+5N(0,1)<55)=P(-2<N(0,1)<2)=2P(0<N(0,1)<2)=2*0.47725=0.9545"

So, among the 1000 children there will be approximately 1000*0.9545=954.5 "\approx" 955 with such weight